Test whether the series ∞Σn=0 1/(n^5+x^3) converges uniformly or not

1
Expert's answer
2022-03-08T11:55:01-0500

We test the uniform convergence of series only for "x>0". For "x<0" equations "n^5+x^3=0" have solutions for "n\in\mathbb{N}" and therefore some terms under the sum will be equal to infinity. The terms "\frac{1}{n^5+x^3}" can be bounded in the following way: "|\frac{1}{n^5+x^3}|<\frac{1}{n^5}". Series "\sum_{1}^{+\infty}\frac{1}{n^5}" converge due to the integral criterion. Namely, "\int_{1}^{+\infty}\frac{dx}{x^5}=-\frac{1}{4}x^{-4}|_{1}^{+\infty}=\frac{1}{4}" The original series "\sum_{1}^{+\infty}\frac{1}{n^5+x^3}" converge uniformly due to the Weierstrass criterion.